# DataPlatter: Boosting Robotic Manipulation Generalization with Minimal Costly Data

### v1

@@ A.1 Our Dataset
Our dataset are collected in the simulation environment of IsaacSim, which is automatically collected using algorithm similar with Zheng et al. [50]. Each scene is initialized with 4 to 6 objects placed randomly on a table, both in position and orientation. A Franka-Panda 7-DoF robotic arm equipped with a two-finger gripper is initialized with a random end-effector pose. A static camera positioned in front of the table, along with a wrist camera mounted on the gripper, are used to capture the RGB and depth observation of the scene, as shown in Fig. 1(b). During the collection process, a target is sampled from the objects on the table and assigned as the target, and a language instruction is generated using pre-defined templates. At each step, the pose of the gripper, action targets generated by the algorithm, robot joint information, gripper status, images from the cameras, task instruction, and status information of all objects in the scene, are recorded for training and reproducing. We use ray-tracing renderer while generating camera images and evaluation. The target objects used in our experiments are shown in  Fig. 8.
@@
For full-stage trajectories, we first sample collision-free grasp labels on the target, which is densely labeled using the collision model of the object using method of Fang et al. [7]. The agent then performs 6-DoF path planning using CuRobo [39] and executes the generated path. For trajectories that only involve SRP stage, to provide a implementable pipeline in real-world robots, we did not directly read the target information from the simulator. Following the method provided in Fig. 1(c), we first located the target from the RGB image captured by the static camera, after which it is feed to a CNN to detect the target bounding box. With the bounding box, we can acquire the mean depth of the target from the depth image, and calculate its position using the intrinsic and extrinsic of the camera. Then in the approach pose sampling stage, we simply samples an end-effector pose within a range of $10cm$ from the target position, ensuring the gripper is oriented towards the target, after which the paths are planned with spatial occupancy information provided by the depth image, and finally the path is executed by the agent.
A.2 CALVIN Dataset
In our datasets we use 5 tasks in D-D split of CALVIN dataset, and the tasks are shown in Tab. 5. In the dataset, each trajectory is divided into SRP and PIP segments using the methods outlined in Eq. 3. In the experiments presented in Tab. 1, we use a total of 150 trajectories per task, setting the parameter $p_{SRP}=50\%$ for the model trained with our method.
@@ A.2 CALVIN Dataset
Appendix B Experiment Details
B.1 Additional Experiments
@@
Seen Targets ($N_{1}=100,N_{2}=0$) Apple 0.6 0.2 0.75 0.45 0.7 0.5
@@
Novel Targets ($N_{1}=50$) Cleansing Foam 0.65 0 0.35 0.15 0.25 0.25
@@
B.2 Detailed Experiment Results
@@
Object Group Target Object None ($N_{1}=0,N_{2}=0$) Only SRP ($N_{1}=0,N_{2}=200$) Full-Stage ($N_{1}=200,N_{2}=0$)
Seen $N_{1}=200$ $N_{2}=0$ Apple 0.65 0.5 0.5
@@
Object Group Target Object None ($N_{1}=0,N_{2}=0$) Only SRP ($N_{1}=0,N_{2}=100$) Full-Stage ($N_{1}=100,N_{2}=0$)
Seen $N_{1}=100$ $N_{2}=0$ Apple 0.35 0.4 0.7

### v2

@@ B.1 Our Dataset
Our dataset are collected in the simulation environment of IsaacSim, which is automatically collected using algorithm similar with Zheng et al[50]. Each scene is initialized with 4 to 6 objects placed randomly on a table, both in position and orientation. A Franka-Panda 7-DoF robotic arm equipped with a two-finger gripper is initialized with a random end-effector pose. A static camera positioned in front of the table, along with a wrist camera mounted on the gripper, are used to capture the RGB and depth observation of the scene, as shown in Fig. 1(b). During the collection process, a target is sampled from the objects on the table and assigned as the target, and a language instruction is generated using pre-defined templates. At each step, the pose of the gripper, action targets generated by the algorithm, robot joint information, gripper status, images from the cameras, task instruction, and status information of all objects in the scene, are recorded for training and reproducing. We use ray-tracing renderer while generating camera images and evaluation. The target objects used in our experiments are shown in  Fig. 9.
@@
For full-stage trajectories, we first sample collision-free grasp labels on the target, which is densely labeled using the collision model of the object using method of Fang et al[6]. The agent then performs 6-DoF path planning using CuRobo [39] and executes the generated path. For trajectories that only involve SRP stage, to provide a implementable pipeline in real-world robots, we did not directly read the target information from the simulator. Following the method provided in Fig. 1(c), we first located the target from the RGB image captured by the static camera, after which it is feed to a CNN to detect the target bounding box. With the bounding box, we can acquire the mean depth of the target from the depth image, and calculate its position using the intrinsic and extrinsic of the camera. Then in the approach pose sampling stage, we simply samples an end-effector pose within a range of $10cm$ from the target position, ensuring the gripper is oriented towards the target, after which the paths are planned with spatial occupancy information provided by the depth image, and finally the path is executed by the agent.
B.2 CALVIN Dataset
In our datasets we use 5 tasks in D-D split of CALVIN dataset, and the tasks are shown in Tab. 8. In the dataset, each trajectory is divided into SRP and PIP segments using the methods outlined in Sec. 4. In the experiments presented in Tab. 1, we use a total of 150 trajectories per task, setting the parameter $p_{SRP}=50\%$ for the model trained with our method.
@@ B.2 CALVIN Dataset
Appendix C Experiment Details
C.1 Environment Setup
VLA model
Training data
In the simulation environment of IsaacSim, we generated a dataset for object-picking tasks involving target objects of various categories and geometrical shapes. For SRP-only trajectories, to provide a implementable pipeline in real-world robots, we did not directly read the object information from simulation, instead we applied a detection-sampling method provided in Fig. 1(c). Details are shown in Sec. B.1. During trajectory generation, we observe that the SRP-only trajectories are generated 2.5×2.5$\times$2.5 × faster than those using full-stage data, while the length of the full-stage data is only 1.4×1.4$\times$1.4 × that of the SRP data. In real-world data collection this discrepancy can only be even larger. For the CALVIN dataset used in the experiment, trajectory segments are divided based on the distance and direction of the gripper. Specifically, assume the gripper approach direction vector is $\mathbf{a}_{n}$ (as shown in Fig. 8) and the vector from the gripper to the target object is $\mathbf{v}$, before the gripper manipulates the target object, the trajectory segment is labeled as PIP if $\left\|\mathbf{v}\right\|\leq 0.2m$ and $\left\langle\mathbf{a}_{n},\mathbf{v}\right\rangle\leq\frac{\pi}{3}$, and the segment before PIP is labeled as SRP. Details are shown in Sec. B.2.
Evaluation
Training
$p_{SRP}=\frac{N_{2}}{N_{1}+N_{2}}.$ (7)
C.2 Scaling with SRP Proportion
C.3 Additional Experiments
@@
Seen Targets ($N_{1}=100,N_{2}=0$)
Apple 0.6 0.2 0.75 0.45 0.7 0.5
@@
Novel Targets ($N_{1}=50$)
Cleansing Foam 0.65 0 0.35 0.15 0.25 0.25
@@
C.4 Detailed Experiment Results
@@ 20 Tasks
Object
Group
Target Object
None
($N_{1}=0,N_{2}=0$)
Only SRP
($N_{1}=0,N_{2}=200$)
Full-Stage
($N_{1}=200,N_{2}=0$)
Seen
$N_{1}=200$
$N_{2}=0$
Apple 0.65 0.5 0.5
@@
Object
Group
Target Object
None
($N_{1}=0,N_{2}=0$)
Only SRP
($N_{1}=0,N_{2}=100$)
Full-Stage
($N_{1}=100,N_{2}=0$)
Seen
$N_{1}=100$
$N_{2}=0$
Apple 0.35 0.4 0.7
@@
C.5 Failure Analysis
[…]
Figure 11: Statistical analysis of failure reasons in our experiments. Most failure cases are cause by inaccuracies in manipulation poses, manipulation of the wrong objects, robot getting stuck due to obstacles or kinematic limitations, and failures in locating target object.